This problem is from Dr.Rapalje's Intermediate Algebra book, section 1.05, page 56, problem number 7.
Using a "let" statement:
Find two consecutive odd integers such that twice the second plus the first is 121.

Answer by fractalier(6550)

Let x be the first number.
Then let x + 2 be the next consecutive odd integer.
Now x + 2(x + 2) = 121 so that
x + 2x + 4 = 121
3x + 4 = 121
3x = 117
x = 39
x + 2 = 41